Shift Summary for evening shift, 07/19/99

Patricia McBride (mcbride@fnpsph)
Tue, 20 Jul 1999 00:20:48 CDT

Shift Summary for evening shift, 07/19/99

On Shift: Val, Patty

Others present: Peter Shanahan, Rick Kessler, Sydney, Jim Graham

Hours of beam delivered at all intensities: 4.5
Hours of beam delivered at nominal intensity: 4.5
Hours of E832 data (special runs not included): 2.25
Hours of special runs (laser scans, muon runs...): 0.25

Accounting of lost time: Access to repair clock fanout

========================================================================
Runs ended this shift:

Run 13946 - (13 M events) - BEAM_832
Run 13947 - BEAM_832 (no data)
Run 13948 - MUON (short run)
Run 13949 - PED

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

22:20 Val and Patty make an access to turn off CsI low voltage

23:00 Val and Peter inspect/repair clock fanout

========================================================================
Overview of shift:

Several hours of smooth running ended with Linac problems. No beam
for 3.5 hours. When beam returned we switched to muon mode to
do a DC plateau. Plateau run was aborted due to
problems with the clock fanout for the CsI. Plateau will be done later
this week.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13946 235 16:05 Jim frees disk space on ktev4

260 16:40 Val frees disk space on ktev2

17:45 Linac problems -- run ended

13947 18:20 still no beam

18:45 did gas check list; changed blue ice packs;
stopped run

13948 18:50 setup for muon run

21:13 Beam returns, muon run at nominal voltages
DC1, DC2: 2400 V (2200 on windows)
DC3 : 2400 V
DC4 : 2450 V
USER LATCH = 0

21:30 MI-60 House Bypass Loop tripped on 16V Lambda
supply.

107 21:35 beam returns for 1 spill
DC1, DC2: 2550 V (2350 on windows)
DC3, DC4: 2550 V
USER LATCH = 1
DC3 offspill current at 5.4 microamps

108 21:36 beam lost; MI power supply
CSI problems, slow da shows problem with RF
page Rick, Ed and Peter

22:03 beam is back, but we still have problems with RF

22:10 reset chambers to nominal voltages
DC3 offspill current at 3.9 microamps

22:20 Val and Patty go on access to turn off CsI low
voltage and inspect fanout crates

22:30 end muon run;
abandon DC plateau activities for now

23:00 Peter arrives; Val and Peter go on access

23:35 Fuse is blown on clock fanout, repair underway

23:48 Repair complete

13949 23:49 Pedestal Run;
turn on Magnet (pos) to prepare for beam

23:48 Call control room to end access and
setup for data mode


========================================================================
Items to Follow up on Next Shift